# A quadratic estimation for the Kühnel conjecture on embeddings

### v6

@@ 1. Introduction and main result
For a $2k$-manifold $M$ into which $\Delta_{n}^{k}$ can be embedded, the linear in $n$ estimate $\beta_{k}(M)\geqslant\dfrac{n-2k-1}{k+1}$ is proved in [PT19] [^2] See [KS21] for a simpler [KS21, Remark 1.2.e] exposition of this estimation. Observe that the paper [PT19] contains another result mentioned in Remark 1.4.c. after a weaker linear estimate of [GMP+]. So even linear estimates appeared more than 20 years after the Kühnel conjecture 1.3 on embeddings was put forward, and are published in good journals. For 6 years after the linear estimate in [GMP+], and before this paper, no stronger estimates appeared [Ku23, §4]. There is another interesting approach to the Kühnel conjecture 1.3 on embeddings: the papers [Ad18, AP24] jointly claim it (without explicitly stating this in [AP24]). See §6 including specific critical Remarks 6.2, 6.5 and 6.7.
@@ 2. Improvements, idea of proof and corollaries
The following two sets111111 In the dual language of parallelelepipeds (see footnote 4) Lemma 4.1 states that the following two sets are equal: the set of all ordered pairs of vectors in $[3]^{k+1}$ such that the vectors have no equal components; the mod $2$ sum of products $P\times Q$ over all ordered pairs $(P,Q)$ of parallelepipeds whose intersection is $1^{\times k+1}$. are equal:
@@
The proof is by induction on $m_{1}+\ldots+m_{\ell}$. The base follows by the above simple case. Let us prove the inductive step in the case when there is a unit in the union of under-diagonal blocks. Denote by $D$ the given matrix. Arrange the rows and the columns of $D$ lexicographically121212 The pair $(x,a)$ is said to be lexicographically smaller than $(y,b)$ if either $x=y$ and $a<b$, or $x<y$.. Let $(i,a)$ be the lexicographically maximal (i.e. the lowest) row whose intersection with the union of under-diagonal blocks of $D$ is non-zero. Let $(j,b)$ be the lexicographically minimal (i.e. the leftmost) column whose intersection with the row $(i,a)$ is non-zero. Formally, $(i,a)$ is the lexicographically maximal row such that there is
@@ 6. An approach via generalized Grünbaum-Kalai-Sarkaria conjecture
Here we present an interesting alternative approach [Ad18, AP24] to the asymptotic version 1.4.d of the Kühnel conjecture on embeddings. (This approach does not work for almost- or $\mathbb{Z}_{2}$-embeddability, see Remarks 2.1 and 2.2.) This version follows from Conjectures 6.3.b and 6.6 below. Conjecture 6.3.b in its turn follows from Conjecture 6.1 as explained in Remark 6.4.
We write ‘conjecture’ not ‘theorem’ in 6.1, 6.3 and 6.6 because there is a publicly available but so far unanswered criticism of their proofs (see Remarks 6.2, 6.5 and 6.7). These remarks show that the argument in [Ad18, AP24] for Conjectures 6.1, 6.3 and 6.6 is not reliable up to the standards of refereed journals, hence these papers do not present a legitimate claim for the conjectures. So we do not present other critical remarks; we do not certify that the other parts of these papers are reliable. Also observe that these papers are unpublished. Besides, Remark 1.4.e suggests that at least one of Conjectures 6.1 and 6.6 is wrong or hard to prove, so that a (dis)proof of the Kühnel conjecture 1.3 on embeddings would require technique different from the one used for Conjectures 6.1 and 6.6.
We sent the criticism of Remarks 6.2 and 6.5 to K. Adiprasito before putting to arXiv version 4 of the current paper (containing the remarks). We sent the criticism of Remark 6.7 to K. Adiprasito and Z. Patáková before putting to arXiv version 6 of the current paper (containing the remarks). At the time of writing neither we receive any reply nor the corresponding updates of [Ad18, AP24] appeared on arXiv.
@@
(a) Conjecture 6.6 is stated as a result in [AP24, Theorem 2]. For criticism of [AP24] see important remarks (b,c) and minor remarks (d-h) below.
We are grateful to K. Adiprasito, S. Melikhov, P. Patak, B. Sanderson, and M. Tancer, e-mail exchange with whom helped us to conclude that Conjecture 6.6 was open in April 2024 (and so was open in 2022 when the first versions of this paper appeared on arXiv).141414 This contradicts a referee report of March 2024 (on shortened version 2 of this paper). The report wrongly treated Conjecture 6.6 as trivial, and so wrongly claimed that the Kühnel conjecture 1.3 for PL embeddings is solved by [Ad18]. Based only upon this wrong claim, the report recommended rejection. So Conjecture 6.6 remains open until the indicated gaps of [AP24] are filled (or until a proof following another approach is written).
(b) In p. 3, proof of Lemma 6, it is not proved that, and it is not clear why
@@
(c) The introduction of [AP24] hides the authors’ claim for (the asymptotic version 1.4.d of) the Kühnel conjecture on embeddings. (Recall that this version follows from Conjectures 6.3.b and 6.6.) This claim was put forward at the beginning of April 2024 by K. Adiprasito [^15] In order to avoid confusion, here we present our April 14, 2024 letter to K. Adiprasito. Conjectures 1.3.f and 4.1.a of this letter are Conjectures 6.6 and 6.1 of this paper.
@@
, cf. footnote 14. Instead of presenting this claim, the introduction of [AP24] writes: ‘Some people may consider this theorem [Conjecture 6.6 = [AP24, Theorem 2]] as a folklore’. This is misleading because this creates the impression that the proof of Conjecture 6.6 in [AP24] is a minor folklore result, not a claim for a famous conjecture, so the proof need not be carefully checked, and potential critical remarks on the proof only concern minor technicalities. The proof is quite technical, so the critical remarks on the proof are necessarily technical. (For an intuitive remark see Remark 1.4.e.)
@@
(e) In p. 3, proof of Lemma 7, ‘the obvious fact’ is not proved and needs to be proved
(f) In p. 1, Example 3, both pairs $(\sigma,\partial\sigma)$ and $(n*\gamma,\gamma)$ are homeomorphic to $(D^{2},\partial D^{2})$. So the homeomorphism $\varphi$ between these pairs cannot map some interior point of $\sigma$ to $\gamma$. Besides, ‘A flat neighborhood with respect to’ is undefined and is unclear.
@@
(i) There are complexes PL embeddable into $\mathbb{R}^{d}$ but for which there are no embeddings into $\mathbb{R}^{d}$ affine on any simplex [vK41, PW]. See stronger results in [FK13, Theorems 1.4, 2.1, 2.4]. Although these results do not refute Conjecture 6.6 (cf. (d)), it is interesting if the complexes can be used to construct counterexamples to Conjecture 6.6.

### v8

@@ 1. Introduction and main result
For a $2k$-manifold $M$ into which $\Delta_{n}^{k}$ can be embedded, the linear in $n$ estimate $\beta_{k}(M)\geqslant\dfrac{n-2k-1}{k+1}$ is proved in [PT19] [^2] See [KS21] for a simpler [KS21, Remark 1.2.e] exposition of this estimation. Observe that the paper [PT19] contains another result mentioned in Remark 1.4.c. after a weaker linear estimate of [GMP+]. So even linear estimates appeared more than 20 years after the Kühnel conjecture 1.3 on embeddings was put forward, and are published in good journals. For 6 years after the linear estimate in [GMP+], and before the current paper, no stronger estimates appeared [Ku23, §4].
There is another interesting approach to the Kühnel conjecture 1.3 on embeddings: the papers [Ad18, AP24] jointly claim it (without explicitly stating this in [AP24]). In §6 we justify that the realization of this approach in [Ad18, AP24] is not reliable up to the standards of refereed journals, by specific critical Remarks 6.2, 6.5 and 6.7333We welcome public statements of different reliability standards [Sk21d]. So far none have appeared, see Remark 6.11.
Because of a referee report we received, we have to recall that arXiv version 2 of the current paper (where the proof is complete, or at least not criticized) appeared on September 1, 2022, earlier than [AP24].
@@
  • (g)
    Shortened versions of this paper were rejected from Intern. Math. Res. Notices upon incompetent reports, see [Sk21d, Examples 6.7 and 6.8]. This is some evidence that the results of this paper are interesting and reliable.
@@ 2. Improvements, idea of proof and corollaries
The following two sets121212 In the dual language of parallelelepipeds (see footnote 5) Lemma 4.1 states that the following two sets are equal: the set of all ordered pairs of vectors in $[3]^{k+1}$ such that the vectors have no equal components; the mod $2$ sum of products $P\times Q$ over all ordered pairs $(P,Q)$ of parallelepipeds whose intersection is $1^{\times k+1}$. are equal:
@@
The proof is by induction on $m_{1}+\ldots+m_{\ell}$. The base follows by the above simple case. Let us prove the inductive step in the case when there is a unit in the union of under-diagonal blocks. Denote by $D$ the given matrix. Arrange the rows and the columns of $D$ lexicographically131313 The pair $(x,a)$ is said to be lexicographically smaller than $(y,b)$ if either $x=y$ and $a<b$, or $x<y$.. Let $(i,a)$ be the lexicographically maximal (i.e. the lowest) row whose intersection with the union of under-diagonal blocks of $D$ is non-zero. Let $(j,b)$ be the lexicographically minimal (i.e. the leftmost) column whose intersection with the row $(i,a)$ is non-zero. Formally, $(i,a)$ is the lexicographically maximal row such that there is
@@ 6. An approach via generalized Grünbaum-Kalai-Sarkaria conjecture
Here we present an interesting alternative approach [Ad18, AP24] to the asymptotic version 1.4.d of the Kühnel conjecture on embeddings. (This approach does not work for almost- or $\mathbb{Z}_{2}$-embeddability defined in Remarks 2.1 and 2.2.) This version follows from Conjectures 6.3.b and 6.6 below. Conjecture 6.3.b in its turn follows from Conjecture 6.1 as explained in Remark 6.4.
We write ‘conjecture’ not ‘theorem’ in 6.1, 6.3 and 6.6 because there is a publicly available but so far unanswered criticism of their proofs (see Remarks 6.2, 6.5 and 6.7). These remarks show that the argument in [Ad18, AP24] for Conjectures 6.1, 6.3 and 6.6 is not reliable up to the standards of refereed journals, hence these papers do not present legitimate claims for the conjectures. So we do not present other critical remarks; we do not certify that the other parts of these papers are reliable. Besides, Remark 1.4.e suggests that at least one of Conjectures 6.1 and 6.6 is wrong or hard to prove, so that a (dis)proof of the Kühnel conjecture 1.3 on embeddings would require technique different from the one used for Conjectures 6.1 and 6.6141414The paper [Ad18] is unpublished. Historical note. The paper [AP24] appeared on arXiv in April 2024, after arXiv version 2 of the current paper (where the proof is complete, or at least not criticized) appeared in September, 2022. So the paper [AP24] was apparently submitted to a journal after the submission (of the shortened arXiv version 3) of the current paper to a journal in February, 2024. See also footnotes 16 and 18..
Suppression of criticism. We sent to K. Adiprasito the criticism of [Ad18] from Remarks 6.2 and 6.5 in April 2024, before putting to arXiv version 4 of the current paper (containing the remarks). So far neither we receive any reply nor an update of [Ad18] corrected according to the criticism, or referring to the criticism, appeared on arXiv. We sent to K. Adiprasito and Z. Patáková the criticism of (arXiv version 2 of) [AP24] from Remarks 6.7.(b–h) in February 2025 before putting to arXiv version 6 of the current paper (containing the remark). After the appearance of this version 6 on arXiv, we received letters expressing disagreement with the criticism of [AP24], but not justifying (and even not mentioning) disagreement with any of the specific Remarks 6.7.(b–h); see our replies in Remark 6.11. In (arXiv version 3 and journal version of) [AP24] all the specific remarks (except a part of minor Remark 6.7.f from arXiv version 6 of the current paper) are ignored, and no reference to the criticism is presented. We sent to the Editors of Bull. LMS the criticism of (published version of) [AP24] from Remarks 6.7.(b–i) in March 2025. The Editors’ reply (Remark 6.8), its criticism (Remark 6.10), lack of the Editors’ reply to the criticism, and no information on the criticism on the journal’s site, could be interesting to math community, because they show that the reliability standards of Bull. LMS are surprisingly low.
@@
(a) Conjecture 6.6 is stated as a result in [AP24, Theorem 2]. For criticism of [AP24] see important specific remarks (b,c,i) and minor specific remarks (d–h) below. In all these remarks (b–h) we refer to abstract, introduction, lemmas, etc. from arXiv versions 2 and 3 of [AP24] (all the remarks are relevant to the published version).
We are grateful to K. Adiprasito, S. Melikhov, P. Patak, B. Sanderson, and M. Tancer, e-mail exchange with whom helped us to conclude that Conjecture 6.6 was open in April 2024 (and so was open in 2022 when the first versions of the current paper appeared on arXiv).161616 This contradicts a referee report of March 2024 (on the shortened arXiv version 3 of the current paper). The report wrongly treated Conjecture 6.6 as trivial, and so wrongly claimed that the Kühnel conjecture 1.3 for PL embeddings is solved by [Ad18]. Based only upon this wrong claim, the report recommended rejection. So Conjecture 6.6 remains open until the indicated gaps of [AP24] are filled (or until a proof following another approach is written).
The fact that the paper [AP24] is published does not change the situation. The paper was not properly refereed, because the referee did not give even those specific critical remarks which are obvious (even to a non-specialist in the area), see e.g. (d). Now that the responsibility for the claim is shared by the journal, we sent our critical remarks to the Editors (see Remark 6.12). Having their answer (or having none) would expose the reliability standards of the journal.
(b) In p. 3, proof of Lemma 6, it is not proved that, and it is not clear why171717 (added in this arXiv version 8 comparatively to version 6) We are grateful to a clever mathematician (who expressed his wish to remain anonymous) for his efforts to prove the statements highlighted in (b), and to the Editors of Bull. LMS for their public reply to (b) (see Remarks 6.8.b and 6.10.b). These efforts and reply helped us to conclude that leaving the statements highlighted in (b) without proof is not up to the publicly asserted reliability standards of refereed journals, cf. [Sk21d].
@@
(c) The introduction hides the claim for (the asymptotic version 1.4.d of) the Kühnel conjecture on embeddings. (Recall that this version follows from Conjectures 6.3.b and 6.6.) This claim was put forward at the beginning of April 2024 by K. Adiprasito181818 In order to avoid confusion, here we present our April 14, 2024 letter to K. Adiprasito. Conjectures 1.3.f and 4.1.a of this letter are Conjectures 6.6 and 6.1 of the current paper.
@@
, cf. footnote 16. Instead of presenting this claim, the introduction reads: ‘Some people may consider this theorem [Theorem 2] as a folklore’. This is misleading because this creates the impression that the proof of Conjecture 6.6 in [AP24] is a minor folklore result, not a claim for a famous conjecture, so the proof need not be carefully checked, and potential critical remarks on the proof only concern minor technicalities. The proof is quite technical, so the critical remarks on the proof are necessarily technical. (For an intuitive remark see Remark 1.4.e.)
@@
(e) In p. 3, proof of Lemma 7, ‘the obvious fact’ is not proved and needs to be proved.
(f) (updated in this arXiv version 7 comparatively to version 6) In p. 1, Example 3, ‘A flat neighborhood with respect to’ is undefined and is unclear; the explanation ‘the embedding is homeomorphic to the standard embedding’ is meaningless because a homeomorphism a relation between spaces not maps.
@@
00111122223333444455556666
0202020211113333444455556666
Figure 1. The (red) complex $K=(0-2-4)$ in the triangulation $M$ (left). The (red) complex $K^{\prime}=(02-4)$ in the triangulation $M^{\prime}$ (right). Both $K^{\prime},M^{\prime}$ are obtained from $K,M$ by valid contraction of the edge $(0-2)$. The complex $K$ is obtained from $K^{\prime}$ by edge subdivision, but $M$ cannot be obtained from $M^{\prime}$ by edge subdivision.
(i) (added in this arXiv version 7 comparatively to version 6) P. 4, proof of Theorem 2. The proof does not explicitly use the assumption that $M$ is a PL manifold, although the analogue of Theorem 2 for complexes that are not manifolds is incorrect. Presumably, the usage of this assumption is hidden in the phrase ‘It follows that the resulting triangulation $M^{\prime}$ of $M$ is the desired one’. In this phrase it is not proved that the obtained complex $M^{\prime}$ is PL homeomorphic to $M$, i.e. is indeed a triangulation of $M$. This could be proved using the following conjecture (whose analogue for complexes $K$ is incorrect): the complex obtained from a triangulation $K$ of a PL manifold by valid edge contraction is PL homeomorphic to $K$. The paper gives no proof or reference to this conjecture [^19] Even if this is only lack of an explicit statement and a reference, not a non-trivial conjecture used without an explicit statement and a proof, this contributes to gradual loss of knowledge of PL topology, of which the authors complain before Example 3.
For a reader’s convenience recall the used result of Alexander [2, Corollary 10:2d and page 299] and Newman [6, pages 187–188] (we use the citation as in [AP24]): any two of PL homeomorphic complexes can be obtained one from the other by a sequence of edge subdivisions and inverse operations, cf. [Go13, page 2].
(and no other proof that $M^{\prime}$ is PL homeomorphic to $M$).
The conjecture is non-trivial because there is a valid edge contraction $\Psi$ in a triangulation of a PL manifold such that $\Psi$ is not an inverse to edge subdivision, although the restriction of $\Psi$ to some subcomplex is an inverse to edge subdivision in the subcomplex (Figure 1).
@@
There are complexes PL embeddable into $\mathbb{R}^{d}$ but for which there are no embeddings into $\mathbb{R}^{d}$ affine on any simplex [vK41, PW]. See stronger results in [FK13, Theorems 1.4, 2.1, 2.4]. Although these results do not refute Conjecture 6.6 (cf. Remark 6.7.d), it is interesting if the complexes can be used to construct counterexamples to Conjecture 6.6.
Remark 6.8.
Below we present the Editors’ reply to the criticism of Remark 6.7 (technically speaking, to the letter of Remark 6.12.a). The Editors kindly allowed to publish their reply (cf. Remark 6.12.b). This cannot but deserve respect. However, we are sorry to justify in Remark 6.10 that the reply is not up to the reliability and clarity standards expected from a refereed journal (see Example 6.9). References in the reply are to [AP24], except the place where we added ‘[DS: of arXiv:2208.04188v7]’.
Apr 12, 2025Dear Professors Skopenkov and Dzhenzher,
Many thanks for your patience. Following receipt of your message dated 31 March concerning the paper ’A higher dimensional version of Fáry’s theorem’ [Bull. London Math. Soc., February 2025], we consulted the Section Editor for Geometry and Topology, the Handling Editor of the paper, and the expert referee. They do not concur with your assessment that there are important gaps in the proof of Theorem 2 of the aforementioned article.
Specifically, regarding the mathematical portions of Remark 6.7 in arXiv:2208.04188v7 (note that it is this version we have examined), we have agreed upon the following responses:
(b) The first bullet point follows from the description of biased barycentric subdivision (sometimes called partial barycentric subdivision) in terms of chains given in the proof of the lemma, and illustrated in Figure 1.2, while the second bullet point follows from the definition of induced subcomplex (usually called full subcomplex).
(c) This paper itself is making no direct claims about the Kuehnel conjecture. Separately, the question that is addressed in the main result of the paper does not appear to be a ‘famous conjecture’.
(d) It is indeed a simplicial rather than an affine version, but nevertheless very much a Fáry-type theorem. As such, the title (and the other phrases referred to) are acceptable (as a certain level of critical engagement can be expected of the reader).
(e) The ‘obvious fact’ follows from the definition of ’strongly induced subcomplex’ (though some further details may indeed have helped the reader).
(i) "collapsing a valid edge in a PL manifold is a PL homeomorphism” is an elementary fact in PL topology, not a conjecture. (The fact that "not every edge contraction is the inverse of an edge subdivision” does not change this.)
Regarding the paragraph preceding Remark 6.8 [DS: of arXiv:2208.04188v7], there is no reason to believe that the examples of complexes that embed PL but not affinely in $\mathbb{R}^{d}$ would lead to counterexamples to Theorem 2.
Overall, we have found the exposition of the paper to be within the bounds of clarity and correctness that we normally expect of publications in the journal.
In this instance, it is our collective view that the authors of [AP25] do not have a case to answer. We regard the matter as closed.
Sincerely,
Minhyong Kim and Julia Wolf
Example 6.9 (on standards of reply to a public criticism, cf. [Sk21d]).
The reliability standards expected from refereed journals require the following.
A reply to a public requirement (from a scientist having some decent papers in the area, e.g. in PL topology) of a reference to a statement should give either a reference to (a specific result in) a textbook, or a several-line deduction from (specific results in) a textbook.
A reply to a public criticism of a proof by a scientist (having some decent papers in the area) should assume that the scientist have read the proof. So the scientist’s critial remark ‘it is not proved that, and it is not clear why [a statement]’ means ‘it is not proved that, and it is not clear why [a statement], even taking into account the definitions of the objects involved in the statement, and the previous sentences of the proof’. [^20] The burden of proof lies with those who assert the statement.
By public requirement of details of the proof (in particular, of a reference) a scientist takes the reputation risk of missing a trivial proof (in particular, a reference in a textbook). So it is only natural that those who assert the statement take the risk of serving scientific community, either by providing necessary details of the proof, or by seeing that the missing argument they have in mind is fallible (in particular, that the necessary result is not present in textbooks).
See e.g. post-publication arxiv updates of [Sk05, Sk08, GS06] (prepared after private not public criticism).
In particular, nothing of the following can contribute to a reliable justification of a statement.
(a) The information that the statement follows from
$\bullet$ the definitions of the objects involved in the statement, or
$\bullet$ the previous argument, even if the relevant two of the four sentences of the argument are specified.
See e.g. Remark 6.8.b.
(b) A figure, especially a low-dimensional figure in the proof of a higher-dimensional statement. See e.g. Remark 6.8.b.
(c) An information on different terminology. See e.g. Remark 6.8.b.
(d) Naming the statement elementary, without giving either a reference or a deduction. (If this comes from a person or from an organization of authority, then this a logical fallacy ‘argument from authority’ [AA].) See e.g. Remark 6.8.i.
(e) A speculation that the non-triviality of the statement is not sufficiently illustrated by some given fact. See e.g. Remark 6.8.i, the sentence in parenthesis.
(f) Lack of a counterexample to the statement (or lack of reasons to believe that some ideas to construct a counterexample work). See e.g. Remark 6.8, after (i).
(g) A ‘straw man’ logical fallacy [SM]. See e.g. Remark 6.10.cj.
If a reply to the above public requirement / critisism states that the authors of a criticized text do not have a case to answer, then the replying authors / organization should publish a reference to the criticism. This would show that the statement ‘do not have a case to answer’ stands the test of being read together with the critisism. Lack of publication of a reference to the criticism is suppression of criticism. Such lack of publication shows that the authors of a criticized text do have a case to answer, but attempt to hide the fact that their reliability standards are lower than those expected from refereed journals.
Remark 6.10.
(a) In (b,c,i,j) we justify that the reply from Remarks 6.8.bci (and from the sentence after (i)) to the critical Remark 6.7 is not up to the reliability and clarity standards described in Example 6.9. Thus the (important) critical Remarks 6.7.bi remain unanswered.
In (c,d) we show that the clarity standards (see [Sk21d, footnote 1]) of Bull. LMS leave it to the readers to cope with as misleading title and introduction as
$\bullet$ hiding the claim for a famous conjecture; [^21] The result of such a hiding is distraction of people’s attention from reliability of the proof, and from unanswered public criticism of the proof, cf. Remark 6.2.
$\bullet$ giving name ‘a higher-dimensional version of theorem A’ to a result which is different from theorem A in low dimensions.
In the items below, the necessary material from Remark 6.7 is shortly recalled (so the justification can be read without reading Remark 6.7, unless a reader wants to check that our references to Remark 6.7 are faithful).
(b) Remark 6.7.b asked for a proof of two statements (bullet points) in the proof of [AP24, Lemma 6]. The reply (Remark 6.8.b) consists only of
$\bullet$ the information that the first statement follows from the previous two sentences (of the previous four sentences) in the proof of [AP24, Lemma 6]: ‘The first bullet point follows from the description of biased barycentric subdivision in terms of chains given in the proof of the lemma…’;
$\bullet$ the information that ‘the second bullet point follows from the definition of induced subcomplex’ (that is, from the definition of the objects involved in the second bullet point);
$\bullet$ an information on different terminology: ‘(sometimes called partial barycentric subdivision)’, ‘(usually called full subcomplex)’;
$\bullet$ a reference to a low-dimensional figure: ‘and illustrated in Figure 1.2’.
See Example 6.9.abc. [^22] Not only we have carefully read the proof, but we discussed possibility of recovery of the unproved statements with mathematicians of different nationalities and backgrounds, see footnote 17.
(c) Remark 6.7.c states and justifies that ‘The introduction hides the claim for … the Kühnel conjecture on embeddings’. The reply (Remark 6.8.c) ignores the raised problem of ‘hiding’. In particular, the phrase ‘the question that is addressed in the main result of the paper does not appear to be a ‘famous conjecture’’ is the ‘straw man’ logical fallacy [SM] (indeed, the critical Remark 6.7.c does not assert the opposite, i.e that ‘the question that is addressed in the main result of the paper is a ‘famous conjecture”).
(d) The reply (Remark 6.8.d) agrees with the critical Remark 6.7.d (although the reply uses different wording).
(i) Remark 6.7.i asked for a reference or a proof of a statement used (without being explicitly stated) in the proof of [AP24, Theorem 2]. The reply (Remark 6.8.i) does not provide a reference or a proof, but only states that the statement ‘is an elementary fact in PL topology’. (The parenthetical remark in the reply uses this unjustified statement as correct but does not contribute to its justification.). See Example 6.9.de. [^23] See https://old.mccme.ru//circles//oim/home/LIBRARY.pdf. We certainly read [RS72], and before asking for the reference consulted it once again.
(j) The paragraph preceding Remark 6.8 is not a part of critical Remark 6.7. So its discussion does not contribute to answering the criticism, and its inclusion into a reply to critical Remark 6.7 is the ‘straw man’ logical fallacy [SM]. See Example 6.9.fg.
Conventions for the following remarks. In the letters the references are updated. In spite of that, the references are given to the very version of a paper cited in the letter. Otherwise the letters are not changed (the grammar is not corrected).
Remark 6.11.
In order to avoid confusion, we present here our letter to the authors of [AP24], and our replies to the letters from J. D. Paik and K. Adiprasito on our criticism of (arXiv version 2 of) [AP24]. (They did not give us their permission to publish their letters.)
(February 25, 2025)  Dear Zuzana, Dear Karim,
Attached please find a planned update of arXiv:2208.04188 [added later: this update is arXiv:2208.04188v6]. We would be grateful for any comments or remarks, but please do not feel obliged.
Remark 6.7 contains some criticism of your interesting paper arXiv:2404.12265v2. It would be nice if you could update your paper on arXiv according to this criticism.
In order to make this discussion responsible, let us make it public. We perpetually release copyright for this letter to the public domain. If you kindly choose to reply to this letter, not just update your paper on arXiv, could you please perpetually release copyright for your reply to the public domain.
Sincerely Yours, Slava and Arkadiy.
(March 8, 2025)  Dear Joshua David,
Thank you for your letter disagreeing with our (arXiv:2208.04188v6) criticism of an unjustified claim to a ‘new triangulation version’ of Fáry’s theorem (Conjecture 6.6 from arXiv:2208.04188v6), and mentioning Dima Burago, Yuri Burago, and Anton Petrunin. Our criticism is justified by specific critical remarks of Remark 6.7 from arXiv:2208.04188v6. Your letter does not disagree with any of those specific critical remarks. So in our opinion, your letter jeopardizes your reputation as a scientist.
If your opinion is different, we welcome its responsible pronunciation. In order to make this discussion responsible, we have to make it public. We encourage you to put your disagreement to arXiv, or to perpetually release copyright for your letter to the public domain. We would be glad to present critical remarks to (or praise) any text publicly available on arXiv (or in the public domain by author’s perpetually releasing copyright), and relevant to our paper.
Sincerely Yours, Slava and Arkadiy.
(March 8, 2025)  Karim,
This is libel, so please make it public so that you would assume responsibility for the libel.
A.
(March 9, 2025)  Dear Karim,
Thank you for forwarding us your letter to Josh naming our criticism (arXiv:2208.04188v6, of an unjustified claim to a ‘new triangulation version’ of Fary’s theorem) ‘incompetence’ . Our criticism is based on specific critical remarks of Remark 6.7 of arXiv:2208.04188v6. Your letter contains no criticism of those specific critical remarks. So, in our opinion, your letter jeopardizes your reputation as a scientist.
If your opinion is different, we welcome its responsible pronunciation. In order to make this discussion responsible, we have to make it public. We encourage you to put your disagreement to arXiv, or to perpetually release copyright for your letter to the public domain. We would be glad to present critical remarks to (or praise) any text publicly available on arXiv (or in public domain by author’s perpetually releasing copyright for the text), and relevant to our paper.
Sincerely Yours, Slava and Arkadiy.
Remark 6.12.
In order to avoid confusion, we present here our letters to the Editors of the journal where the paper [AP24] was published.
(a) March 31, 2025
Dear Managing Editors Minhyong Kim and Julia Wolf, Dear Editors,
Hope you are fine and healthy.
We write this letter on the understanding that LMS strives to ensure that its journals’ contribution to the published record is reliable (https://www.lms.ac.uk/publications/policies/ethicalpolicy). We feel obliged to inform you that Theorem 2 claimed in
is not reliable, and the introduction is misleading. See justification by specific critical Remarks 6.7 (b,c,d,e,g,h) of [DS] := arXiv:2208.04188v6. In these remarks we refer to the arXiv:2404.12265v3 version of [AP24]; all the remarks are relevant to the published version. See also Remark 6.7.i from the attached update of [DS] (added later: this update is arXiv:2208.04188v7).
Now that the responsibility for the claim is shared by the journal, it would be nice if the Editors could ask for the authors’ and the referees’ answers to these specific remarks. Moderation of such a discussion by the Editors (or just giving reference to arXiv published criticism of the paper published by LMS), would confirm high reliability and clarity standards of the journal. The result of such a discussion could be either rejection of the paper, or reliable revision of the paper, or arXiv publication of answers to our (arXiv published) specific critical remarks. In order not to mislead the readership, either the publication could be suspended, or a reference to the ongoing discussion of reliability of the proof (and clarity of the introduction) could be added. We would only be glad if the revision taking into account [DS, Remarks 6.7 (b,c,d,e,g,h)] (and Remark 6.7.i from the attached update of [DS]) could be published.
Sincerely Yours, Arkadiy Skopenkov and Slava Dzhenzher.
PS We sent the criticism to the authors of [AP24] even before the v6 update of [DS]. After the appearance of the v6 update, we received letters expressing disagreement with our criticism. Although our criticism is justified by specific critical remarks, disagreement with our criticism did not consider any of those specific critical remarks. In order to avoid confusion, we present our replies to the disagreement in Remark 6.8 from the attached update of [DS]. Remark 6.9 of that update is a copy of this letter. We plan to publish the information on the Editors’ reply (or on the lack of it).
PSS We do understand that the paper [AP24] was refereed. However, the referee did not give even those specific critical remarks which are obvious (even to a non-specialist in the area), see e.g. [DS, Remark 6.7.d]. So it is not surprising that a referee missed a gap in the (quite technical) proof.
(b) April 13, 2025
Dear Minhyong Kim and Julia Wolf,
Many thanks for your reply to our critical remarks on the paper
[AP25] := [Adiprasito-Patakova, Bull. London Math. Soc., February 2025].
The reply is very interesting to the math community, as a revelation of the bounds of clarity and correctness that the Editors of Bull. LMS journals normally expect of publications in the journal. So would you kindly allow us to publish the reply? Technically speaking, would you release the copyright for the reply to the public domain? Please feel free to make any changes in your reply before making it public.
In an update of arXiv:2208.04188 we shall either add your reply, or this letter and the information that you did not allow us to make the reply public. The first option would be a sound confirmation that our critical remarks on [AP25] are so far answered in a responsible way. We are afraid that the second option would be a sound confirmation that our (openly published in arXiv:2208.04188) critical remarks on [AP25] are justified, and cannot be answered in a responsible way.
Sincerely Yours, Arkadiy Skopenkov and Slava Dzhenzher.
(c) June 19, 2025
Dear Managing Editors Minhyong Kim and Julia Wolf,
Hope you are fine and healthy.
Thank you for allowing us to publish your reply to a criticism of a paper published in Bull. LMS,
This cannot but deserve respect. However, we are sorry to justify (in Remark 6.10) that the reply (Remark 6.8) is not up to the reliability and clarity standards expected from a refereed journal (see Example 6.9). The numbering is from the attached update of arXiv:2208.04188. Unless we receive your public reply, we plan to make the update publicly available in a week. Your public reply to Example 6.9 and Remark 6.10 is very interesting to the math community, as a revelation of the bounds of clarity and correctness that the Editors of Bull. LMS journals normally expect of publications in the journal.
We value our collaboration with the London Mathematical Society very much (the more so because Arkadiy is grateful to LMS for the 2002 invitation to deliver research lectures in Great Britain). More specifically, we would be pleased
$\bullet$ to postpone arXiv update of arXiv:2208.04188, and
$\bullet$ to publish in a later revised update, instead of your reply (Remark 6.8) and its criticism (Remark 6.10), a revised version of your reply (which hopefully would not require such a harsh criticism).
For this one needs LMS to publish on the journal’s website the information on the ongoing discussion whether the reliability (and clarity) standards of [AP24] are up to the standards of LMS, with a link to the published criticism of [AP24] in §6 of arXiv:2208.04188v7.
This would be a revelation of the high transparency standard of LMS (see arXiv:2101.03745v3, Remark 4.2.a), and its respect to responsible professional criticism.
Sincerely, Arkadiy Skopenkov and Slava Dzhenzher.